# Phase diagram of the interacting Haldane model with spin-dependent sublattice potentials

### v1

@@ Abstract
Using the exact-diagonalization (ED) and mean-field (MF) approaches, we study the ground-state phase diagram of the interacting Haldane model on the honeycomb lattice, taking into consideration of the spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$. Here $\alpha=\text{A}$,B and $\sigma=\uparrow$$\downarrow$ indicate the sublattice and spin components, respectively. If we set $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$), the system favors a spin ordered state. On the other hand, introducing the nearest-neighbor Coulomb interaction $V$ can drive the system to be charge ordered. Due to their competition, we find that in both ED and MF approaches, an exotic state with Chern number $C=1$ survives from two locally ordered phases and a topologically ordered phase with $C=2$. In ED method, other properties, including the fidelity metric, the excitation gap and the structure factors, are also used to determine the critical points. In MF method with large enough lattice size, we define the local order parameters and band gap to characterize the phase transitions. The interacting Haldane model and the spin-dependent lattice potential can be realized in an ultracold atom gas, which may serve as a way to detect this intriguing state.
@@ I Introduction
Different from the traditional framework of the Landau-Ginzburg theory that relies on locally defined order parameters resulting from broken symmetries, topological phases have been identified and characterized based on their global, nonlocal propertiesHasan and Kane (2010); Qi and Zhang (2011). Over the past few years, the categorization of topologically ordered states in non-interacting systems has been completed, in terms of different symmetriesChiu et al. (2016); Zhang et al. (2019); Kruthoff et al. (2017); Vergniory et al. (2019); Tang et al. (2019); Schnyder et al. (2008).
@@
On the other hand, interacting topological insulators, which are known as systems encompassing the interplay of topological properties and electronic correlations, have been extensively studied in recent years Rachel (2018); Hohenadler and Assaad (2013). Correlation effects are expected to generate exotic states in the presence of topologically nontrivial conditions, such as the antiferromagnetic topological state in the Bernevig-Hughes-Zhang model with the on-site Hubbard interactionMiyakoshi and Ohta (2013); Yoshida et al. (2013) and the topologically non-trivial phase with $C=1$ ($C$ denotes the Chern number) in the spinful Haldane-Hubbard model on honeycomb latticeHe et al. (2011); Zhu et al. (2014); Wu et al. (2015); Vanhala et al. (2016); Tupitsyn and Prokof’ev (2019); Mertz et al. (2019); Yuan et al. (2023). The origin of this $C=1$ phase is attributed to a spontaneous SU(2) symmetry breaking with one of the spin components in the Hall state and the other one in a localized state. Similar studies include the antiferromagnetic Chern insulator in Kane-Mele-Hubbard model Jiang et al. (2018) and $C=1$ phase in interacting topological models on the square lattice Wang and Qi (2019); Ebrahimkhas et al. (2021). Other attempts like introducing the double exchange processes to the Haldane Hamiltonian can also break the spin symmetry Tran and Tran (2022), while interplays between topology, on-site and nearest-neighbor interactions do not exhibit such phenomenon Wang and Qi (2019); Shao et al. (2021). In experiment, observing these states are still challenging within the field of material science, and thus the trapped cold atoms in controlled optical lattice may provide an alternative and promising way for achieving this purpose. So far, the experimental realization of the topological Haldane model has been reportedJotzu et al. (2014) and the quantum simulations of strongly correlated systems (including the Fermi-Hubbard model) in ultracold Fermi gases can be reviewed in Ref. Hofstetter and Qin (2018); Esslinger (2010); Tarruell and Sanchez-Palencia (2018).
@@
In this paper, inspired by the studies of spin-dependent optical latticeJotzu et al. (2015); Förster et al. (2009); Mandel et al. (2003), we propose an interacting Haldane model on honeycomb lattice with spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$, where $\sigma=\uparrow,\downarrow$ and $\alpha=$A,B represents the spin and sublattice components, respectively. If one set $\Delta_{\uparrow,\text{A}}=+\Delta$, $\Delta_{\uparrow,\text{B}}$$=-\Delta$, $\Delta_{\downarrow,\text{A}}=-\Delta$ and $\Delta_{\downarrow,\text{B}}$$=+\Delta$, the system favors a spin-density-wave (SDW) state. Meanwhile, we introduce the nearest-neighbor Coulomb interaction $V$ to drive the system into a charge-density-wave (CDW) state. Due to their competitions, an intermediate state can be expected and we thus study its ground-state phase diagram. We find that except for the topologically non-trivial phase with Chern number $C=2$ and two topologically trivial phases (SDW and CDW) with $C=0$, an newly generated topological phase with $C=1$ can be observed in both exact-diagonalization (ED) and mean-field (MF) method. Our finding provide a new version of competitions among topology, electronic correlation and lattice potential to realize the exotic state.
@@
The presentation is organized as follows: we introduce the model, methods and relevant quantities in Sec. II. Sections III presents our results based on the ED and MF approaches. Lastly, a conclusion is given in Sec. IV.
@@ II Model and measurements
$\hat{H}_{l}$ $=$ $U\sum_{i}\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}+V\sum_{\langle i,j\rangle,\sigma,\sigma^{\prime}}\hat{n}_{i,\sigma}\hat{n}_{i,\sigma^{\prime}}$ (3)
$+$ $\sum_{i,\sigma}\Delta_{\sigma,\alpha}\hat{n}_{i,\sigma}.$
In Eq. (2), $\hat{c}^{\dagger}_{i,\sigma}$ ($\hat{c}_{i,\sigma}$) is the creation (annihilation) operator for an electron at site $i$ with spin $\sigma=\uparrow{\rm or}\downarrow$. $t_{1}$ ($t_{2}$) is the nearest-neighbor (next-nearest-neighbor) hopping constant and the Haldane phase $\phi_{i,j}=\phi$ ($-\phi$) in the clockwise (anticlockwise) loop is introduced to the next-nearest-neighbor hopping terms. In Eq. (3), $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ is the number operator; $U$ and $V$ are the on-site and nearest-neighbor Coulomb interaction strength, respectively. $\Delta_{\sigma,\alpha}$ is the spin-dependent lattice potential: for spin $\sigma=\uparrow$, $\Delta_{\sigma,A}=+\Delta$ and $\Delta_{\sigma,B}=-\Delta$; for spin $\sigma=\downarrow$, $\Delta_{\sigma,A}=-\Delta$ and $\Delta_{\sigma,B}=+\Delta$.
In what follows, the model in Eq. (1) are named as the extended Haldane-Hubbard model with spin-dependent lattice potentials. Throughout the paper, we set $t_{1}=1$, $t_{2}=0.2$, $\phi=\pi/2$ and focus on the ground-state phase diagram of this model at half-filling.
@@ II.1 Exact diagonalization in real space
Other properties used to characterize the critical behavior of quantum phase transition include the ground-state fidelity metric $g$, which is defined as Zanardi and Paunković (2006); Campos Venuti and Zanardi (2007); Zanardi et al. (2007)
@@
where x represents the parameters V or $\Delta$, and N is the lattice size. $|\Psi(x)\rangle$ [$|\Psi(x+\delta x)\rangle$] is the ground state of $\hat{H}(x)$ [$\hat{H}(x+\delta x)$] and we set $\delta x=10^{-3}$. In addition, the SDW and CDW structure factors can be used to characterize the spin and charge ordered insulators, respectively. Their definitions in a staggered fashion can be written as
$S_{\text{SDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}-\hat{n}_{j,\downarrow})\rangle,$
$S_{\text{CDW}}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}+\hat{n}_{j,\downarrow})\rangle,$ (6)
where $\eta=0$ ($\eta=1$) if sites $i$ and $j$ are in the same (different) sublattice.
@@ II.2 Mean-field method in momentum space
We employ a variational mean-field method in momentum space to contrast the results obtained via the ED method, which has been reported in studying the extended Haldane-Hubbard model without lattice potentials Shao et al. (2021). In our case, by introducing the operators $a_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{A}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$ and $b_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{B}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$, the Hamiltonian (1) can be expressed as
@@
where $\psi^{\dagger}_{\textbf{k}}=[a^{\dagger}_{\textbf{k},\uparrow},b^{\dagger}_{\textbf{k},\uparrow},a^{\dagger}_{\textbf{k},\downarrow},b^{\dagger}_{\textbf{k},\downarrow}]$ is the basis for each lattice momentum ${\bf k}$ and
@@
with densities $n_{\sigma}^{a}=\frac{1}{N}\sum_{\textbf{q}}\langle a_{\textbf{q},\sigma}^{{\dagger}}a_{\textbf{q},\sigma}\rangle_{\rm MF}$ and $n_{\sigma}^{b}=\frac{1}{N}\sum_{\textbf{q}}\langle b_{\textbf{q},\sigma}^{{\dagger}}b_{\textbf{q},\sigma}\rangle_{\rm MF}$. The above mean-field equations can be solved self-consistently by making use of the variational mean-field approach. Once the free energy has converged, the SDW and CDW order parameters can be obtained by
$\mathcal{O}_{\text{SDW}}$ $=$ $\left|\frac{1}{2}\left(\langle\vec{S}_{A}\rangle_{\rm MF}-\langle\vec{S}_{B}\rangle_{\rm MF}\right)\right|,$
$\mathcal{O}_{\text{CDW}}$ $=$ $\left|\left(n^{A}_{\uparrow}+n^{A}_{\downarrow})-(n^{B}_{\uparrow}+n^{B}_{\downarrow}\right)\right|.$ (12)
Here $\vec{S}_{i}=\frac{1}{2}\sum_{\alpha\beta}c_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}c_{i\beta}$, and $\vec{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the vector of spin-1/2 Pauli matrices. Meanwhile, we use the discrete formulation in its multiband (non-Abelian) version to compute the Chern number Fukui et al. (2005).
@@ III Results and analysis
III.1 Results of the Exact Diagonalization
@@
In our ED calculations, we choose the $12$A cluster whose reciprocal lattice contains the $\Gamma$, K, ${}^{\prime}$ points, and one pair of M points, see more details in Ref. Shao et al. (2021). It has been shown that containing the high-symmetry K and ${}^{\prime}$ points are critical to study the quantum phase transition in interacting Haldane systems Shao et al. (2021); Varney et al. (2010, 2011). Meanwhile, we employ the periodic boundary condition and make use of translational symmetries to reduce the Hilbert space size. Therefore, discussions with regard to the excited states below are restricted to the momentum subspace where the ground state belongs to.
@@
The ($V$, $\Delta$) phase diagrams are shown in Figs. 1(a), 1(b), 1(c) and 1(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. The phase transition points (black circles) in Fig. 1 are identified by the positions of the peaks of fidelity metric $g$. Notice that in Fig. 1(a) with $U=0.0$, the results of Chern number matches the phase diagram obtained from the fidelity metric very well, where we use the blue, green and red squares to represent the Chern number $C=2$, $C=1$ and $C=0$, respectively. We can observe from Fig. 1 (a) that the Chern insulator (CI) phase with $C=2$ is in a region with small $V$ and $\Delta$, the spontaneous symmetry breaking SDW or CDW phases with $C=0$ dominates the system when $\Delta$ or $V$ is large enough. Interestingly, an intermediate phase with $C=1$ can be observed surrounded by three expected phases. It is different from the phase diagram in Ref. Vanhala et al. (2016), where the $C=1$ phase is sandwiched between a band insulator and Mott insulator. Notice that their charge ordered phase (band insulator) is governed by the ionic potential $\Delta_{\rm AB}$ which is spin-independent, and their spin ordered phase (Mott insulator) is governed by the on-site Hubbard interaction $U$. While in our case, the charge ordered phase (CDW) is governed by the nearest-neighbor interaction $V$ and the spin ordered phase (SDW) is governed by the spin-dependent lattice potential $\Delta_{\sigma,\alpha}$. We speculate the different competition fashions result in the different locations of $C=1$ phase. In addition, we observe that the $C=1$ phase in Fig. 1(a) can exist with very small $V$, while in small $\Delta$ region ($\Delta\leq 0.5$) it vanishes. This is completely opposite to our results of MF method in Sec. III.2, where the $C=1$ phase can exist in small $\Delta$ region rather than small $V$ region, see Fig. 3(a). We suggest that this is an open issue and expect further studies on it.
@@
Figure 1: Phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of fidelity metric $g$ with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$. The blue, green and red squares indicate results of Chern number $C=2$, $C=1$ and $C=0$, respectively. The black dashed lines in (a) and (d) denote the parameters we choose to show more details below.
@@
When $U=1.0$ in Fig. 1(b), however, results of Chern number do not match the phase diagram obtained from the fidelity metric. Similar discrepancies between Chern number and fidelity metric have also been discussed in Ref. Yuan et al. (2023) and they found that increasing $(N_{x},N_{y})$ from (6, 6) to (20, 20) can effectively reduce such discrepancies. Remember that $N_{x}$ and $N_{y}$ come from the discretized method to calculate Chern number, see the relevant discussions in Sec. II. While here we already use $(N_{x}=20,N_{y}=20)$ to calculate the Chern number, and results of $C=1$ (green points) can not even be observed in Figs. 1(c) and 1(d) with $U=2.0$ and $U=3.0$, respectively. We further use $(N_{x}=60,N_{y}=60)$ to check several points in Fig. 1(d) but with no finding of $C=1$. On the other hand, the area of the intermediate phase identified by the fidelity metric decreases but does not vanishes as $U$ increases from $0.0$ to $3.0$. (We must have a statement here that in our ED discussions, the $C=1$ phase is equivalent to the intermediate phase when $U=0.0$, and they are not equivalent when $U$ is finite). The MF results in Sec. III.2 also show the existence of $C=1$ in the case of $U=3.0$. We thus suggest the issue of whether the $C=1$ phase can exist with $U$ increased to 3.0 may still be open. This issue is worth of further study by other methods like Monte Carlo and density matrix renormalization group because in materials $U$ is generally larger than $V$. Back to Fig. 1(a), the situation with $U=0.0$ and finite $V$ may still be realized in ultracold atom systems due to their advances in parameter simulations.
@@
Figure 2: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $V=0.4$ on the left panels and $\Delta=1.4$ on the right panels. The on-site interaction $U=0$ and the parameters are corresponding to the balck dashed lines in Fig. 1(a).

### v3

@@ Abstract
Using the exact-diagonalization (ED) and mean-field (MF) approaches, we investigate the ground-state phase diagram of the interacting Haldane model on the honeycomb lattice, incorporating spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$. Here $\alpha=\text{A}$,B and $\sigma=\uparrow$$\downarrow$ denote the sublattice and spin components, respectively. Setting $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$) results in the system favoring a spin ordered state. Conversely, introducing the nearest-neighbor Coulomb interaction can induce charge ordering in the system. Due to the competition between these factors, we observe that in both ED and MF approaches, an exotic state with Chern number $C=1$ survives amidst two locally ordered phases and a topologically ordered phase with $C=2$. In the ED method, various properties, such as the fidelity metric, the excitation gaps and the structure factors, are employed to identify critical points. In the MF method, using a sufficiently large lattice size, we define the local order parameters and band gaps to characterize the phase transitions. The interacting Haldane model and the spin-dependent lattice potential may be experimentally realized in an ultracold atom gas, providing a potential means to detect this intriguing state.
@@ I Introduction
In contrast to the traditional framework of the Landau-Ginzburg theory, which relies on locally defined order parameters resulting from broken symmetries, topological phases have been identified and characterized based on their global, nonlocal propertiesHasan and Kane (2010); Qi and Zhang (2011). Over the past few years, the categorization of topologically ordered states in non-interacting systems has been completed, considering various symmetriesChiu et al. (2016); Zhang et al. (2019); Kruthoff et al. (2017); Vergniory et al. (2019); Tang et al. (2019); Schnyder et al. (2008).
@@
On the other hand, there has been extensive research on the interacting topological insulators, which involve the interplay between topological properties and electronic correlations Rachel (2018); Hohenadler and Assaad (2013). Correlation effects are expected to give rise to exotic states in the presence of topologically nontrivial conditions. Examples include the antiferromagnetic topological state in the Bernevig-Hughes-Zhang model with the on-site Hubbard interactionMiyakoshi and Ohta (2013); Yoshida et al. (2013) and the topologically non-trivial phase with $C=1$ ($C$ denoting the Chern number) in the spinful Haldane-Hubbard model on honeycomb latticeHe et al. (2011); Zhu et al. (2014); Wu et al. (2015); Vanhala et al. (2016); Tupitsyn and Prokof’ev (2019); Mertz et al. (2019); Yuan et al. (2023); He et al. (2024). The origin of this $C=1$ phase is attributed to a spontaneous SU(2) symmetry breaking, with one spin component in the Hall state and the other in a localized state. Similar investigations include the antiferromagnetic Chern insulator in Kane-Mele-Hubbard model Jiang et al. (2018) and $C=1$ phase in interacting topological models on the square lattice Wang and Qi (2019); Ebrahimkhas et al. (2021). Additionally, introducing disorder to the interacting Haldane model Silva et al. or double exchange processes to the Haldane Hamiltonian Tran and Tran (2022) can also break the spin symmetry. However, interplays between topology, on-site, and nearest-neighbor interactions do not exhibit such phenomenon Wang and Qi (2019); Shao et al. (2021). In experimental settings, observing these states remains challenging in realistic materials. Trapped cold atoms may offer an alternative and promising approach to achieving this purpose. Notably, the experimental realization of the topological Haldane model has been reportedJotzu et al. (2014). Quantum simulations of strongly correlated systems in ultracold Fermi gases, including the Fermi-Hubbard model, have also been reviewed in Ref. Hofstetter and Qin (2018); Esslinger (2010); Tarruell and Sanchez-Palencia (2018).
@@
In this paper, drawing inspiration from studies on spin-dependent optical latticeJotzu et al. (2015); Förster et al. (2009); Mandel et al. (2003), we propose an interacting Haldane model on honeycomb lattice with spin-dependent sublattice potentials $\Delta_{\sigma,\alpha}$, where $\sigma=\uparrow,\downarrow$ and $\alpha=$A,B represent the spin and sublattice indices, respectively. Setting $\Delta_{\sigma,\text{A}}=+\Delta$ ($-\Delta$) and $\Delta_{\sigma,\text{B}}$$=-\Delta$ ($+\Delta$) for $\sigma=\uparrow$ ($\downarrow$) leads to the system favoring a staggered spin order (SSO) state. While introducing the nearest-neighbor Coulomb interaction $V$ drives the system into a staggered charge order (SCO) state. Due to their competitions, an intermediate state is expected and we explore the ground-state phase diagram of this model. Our findings reveal that, in addition to the topologically nontrivial phase with Chern number $C=2$ and two topologically trivial phases (SSO and SCO) with $C=0$, an newly generated phase with $C=1$ can be observed in both exact-diagonalization (ED) and mean-field (MF) methods. Various properties, including the excitation gaps, the structure factors, and the fidelity metric in ED, as well as the band gap and local order parameters in MF, are obtained to characterize the phase transitions. Especially, an effective potential differences for the two spin species in MF method are defined to analyze the origin of the $C=1$ phase. Our finding presents a perspective on the interplay between topology, electronic correlation and lattice potential, shedding light on the realization of exotic states.
@@
The presentation is structured as follows: we introduce the model, methods and relevant quantities in Sec. II. Sections III presents our results based on the ED and MF approaches. Lastly, a conclusion is provided in Sec. IV.
@@ II Model and measurements
$\hat{H}_{l}$ $=$ $U\sum_{i}(\hat{n}_{i,\uparrow}-\frac{1}{2})(\hat{n}_{i,\downarrow}-\frac{1}{2})$ (3)
$+$ $V\sum_{\langle i,j\rangle}(\hat{n}_{i}-1)(\hat{n}_{j}-1)+\sum_{i,\sigma}\Delta_{\sigma,\alpha}\hat{n}_{i,\sigma}.$
In Eq. (2), $\hat{c}^{\dagger}_{i,\sigma}$ ($\hat{c}^{\phantom{}}_{i,\sigma}$) is the creation (annihilation) operator for an electron at site $i$ with spin $\sigma=\uparrow{\rm or}\downarrow$. $t_{1}$ ($t_{2}$) is the nearest-neighbor (next-nearest-neighbor) hopping constant and the Haldane phase $\phi_{i,j}=\phi$ ($-\phi$) in the clockwise (anticlockwise) loop is introduced to the next-nearest-neighbor hopping terms. In Eq. (3), $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}^{\phantom{}}$ and $\hat{n}_{i}=\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow}$; $U$ and $V$ are the on-site and nearest-neighbor Coulomb interactions, respectively. $\Delta_{\sigma,\alpha}$ is the spin-dependent lattice potential: for spin $\sigma=\uparrow$, $\Delta_{\sigma,A}=+\Delta$ and $\Delta_{\sigma,B}=-\Delta$; for spin $\sigma=\downarrow$, $\Delta_{\sigma,A}=-\Delta$ and $\Delta_{\sigma,B}=+\Delta$.
In what follows, the model in Eq. (1) is named as the extended Haldane-Hubbard model with spin-dependent lattice potentials. Throughout the paper, we set $t_{1}=1$, $t_{2}=0.2$, $\phi=\pi/2$ and focus on the ground-state phase diagram of this model at half-filling.
@@ II.1 Exact diagonalization in real space
Other properties used to characterize the critical behavior include the ground-state fidelity metric $g$, which is defined as Zanardi and Paunković (2006); Campos Venuti and Zanardi (2007); Zanardi et al. (2007)
@@
where x represents the parameters V or $\Delta$, and N is the lattice size. $|\Psi(x)\rangle$ [$|\Psi(x+\delta x)\rangle$] is the ground state of $\hat{H}(x)$ [$\hat{H}(x+\delta x)$] and we set $\delta x=10^{-3}$. In addition, the SSO and SCO structure factors can be used to characterize the spin and charge ordered states, respectively, and their definitions in a staggered fashion can be written as
$S_{\rm SSO}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}-\hat{n}_{j,\downarrow})\rangle,$
$S_{\rm SCO}=\frac{1}{N}\sum\limits_{i,j}{(-1)}^{\eta}\langle(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}+\hat{n}_{j,\downarrow})\rangle,$ (6)
where $\eta=0$ ($\eta=1$) if sites $i$ and $j$ are in the same (different) sublattice, i.e., A or B.
@@ II.2 Mean-field method in momentum space
A variational mean-field method is employed to analyze the ground-state phase diagram of model (1), which has been reported in studying the extended Haldane-Hubbard model without lattice potentials Shao et al. (2021). For our model, by introducing the operators $a_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{A}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$ and $b_{\textbf{k},\sigma}^{{\dagger}}=\frac{1}{\sqrt{N}}\sum_{i\in\text{B}}c_{i,\sigma}^{{\dagger}}e^{\mathrm{i}\textbf{k}\cdot\textbf{r}_{i}}$, the Hamiltonian can be expressed as
@@
where $\psi^{\dagger}_{\textbf{k}}=[a^{\dagger}_{\textbf{k},\uparrow},b^{\dagger}_{\textbf{k},\uparrow},a^{\dagger}_{\textbf{k},\downarrow},b^{\dagger}_{\textbf{k},\downarrow}]$ represents the basis for each lattice momentum ${\bf k}$ and
@@
with densities $n_{\sigma}^{a}=\frac{1}{N}\sum_{\textbf{q}}\langle a_{\textbf{q},\sigma}^{{\dagger}}a_{\textbf{q},\sigma}\rangle_{\rm MF}$ and $n_{\sigma}^{b}=\frac{1}{N}\sum_{\textbf{q}}\langle b_{\textbf{q},\sigma}^{{\dagger}}b_{\textbf{q},\sigma}\rangle_{\rm MF}$. The above mean-field equations can be solved self-consistently by making use of the variational mean-field approach. Once the free energy has converged, the SSO and SCO order parameters can be obtained by
$\mathcal{O}_{\rm SSO}$ $=$ $\left|\frac{1}{2}\left(\langle\vec{S}_{A}\rangle_{\rm MF}-\langle\vec{S}_{B}\rangle_{\rm MF}\right)\right|,$
$\mathcal{O}_{\rm SCO}$ $=$ $\left|\left(n^{A}_{\uparrow}+n^{A}_{\downarrow})-(n^{B}_{\uparrow}+n^{B}_{\downarrow}\right)\right|.$ (12)
Here $\vec{S}_{i}=\frac{1}{2}\sum_{\alpha\beta}c_{i\alpha}^{\dagger}\vec{\sigma}_{\alpha\beta}c^{\phantom{\dagger}}_{i\beta}$, and $\vec{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ is the vector of spin-1/2 Pauli matrices. Meanwhile, we use the discrete formulation in its multiband (non-Abelian) version to compute the Chern number Fukui et al. (2005).
Besides, similar to the definitions in Ref. He et al. (2024); Silva et al. , we define the effective potential differences for spin-$\uparrow$ and spin-$\downarrow$ electrons in the MF method as
$\Delta_{\rm MF}^{\uparrow}$ $=$ $\frac{|\varepsilon_{\uparrow}^{a}-\varepsilon_{\uparrow}^{b}+2\Delta|}{2},$
$\Delta_{\rm MF}^{\downarrow}$ $=$ $\frac{|\varepsilon_{\downarrow}^{a}-\varepsilon_{\downarrow}^{b}-2\Delta|}{2}.$ (13)
Notice that $+2\Delta$ in $\Delta_{\rm MF}^{\uparrow}$ and $-2\Delta$ in $\Delta_{\rm MF}^{\downarrow}$ come from the opposite potential differences for spin $\uparrow$ and $\downarrow$ species.
@@ III Results and analysis
III.1 Results of the exact diagonalization method
@@
In the ED calculations, we choose the $12$A cluster whose reciprocal lattice encompasses the $\Gamma$, K, K points, and one pair of M points, see more details in Ref. Shao et al. (2021). It has been shown that including the high-symmetry K and K points is crucial for studying the quantum phase transition in interacting Haldane models Shao et al. (2021); Varney et al. (2010, 2011). The periodic boundary condition is utilized for the calculation of all properties and when determining the Chern numbers, we make use of translational symmetries to reduce the size of the Hilbert space. The ($V$, $\Delta$) phase diagrams are shown in Figs. 1(a), 1(b), 1(c) and 1(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. In Fig. 1, the phase transition points (black circles) are identified based on the positions of the peaks of fidelity metric $g$.
@@
Let us first focus on the $U=0.0$ case in Fig. 1(a), where the Chern number results match the phase diagram obtained from the fidelity metric very well. Here we use the blue, green and red squares to represent the Chern number $C=2$, $C=1$ and $C=0$, respectively. We can observe from Fig. 1 (a) that the Chern insulator phase with $C=2$ is present in a region with small $V$ and $\Delta$, while the SSO (SCO) phase with $C=0$ dominates the system when $\Delta$ ($V$) is sufficiently large. Interestingly, an intermediate phase with $C=1$ can be observed surrounded by the three expected phases. This is different from the phase diagram in Ref. Vanhala et al. (2016), where the $C=1$ phase is sandwiched between a band insulator and Mott insulator. Notice that their charge ordered phase (band insulator) is governed by the ionic potential $\Delta_{\rm AB}$ which is not spin-dependent, and their spin ordered phase (Mott insulator) is governed by the on-site Hubbard interaction $U$. In contrast, in our case, the staggered charge order (SCO) phase is governed by the nearest-neighbor interaction $V$ and the staggered spin order (SCO) phase is governed by the spin-dependent lattice potential $\Delta_{\sigma,\alpha}$. We propose that the different competition mechanisms result in the different locations of the $C=1$ phase.
@@
Figure 1: Phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of fidelity metric $g$ with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=3.0$. The blue, green and red squares indicate results of Chern number $C=2$, $C=1$ and $C=0$, respectively. The black dashed lines in (a) and (d) denote the parameters we choose to show more details below.
@@
The impact of the on-site interaction $U$ on the phase diagram is depicted in Figs. 1(b), 1(c) and 1(d). In Fig. 1(b) with $U=1.0$, there is a discrepancy between the results of the Chern number and the phase diagram obtained from the fidelity metric. Moreover, results of $C=1$ (green points) can not even be observed in Figs. 1(c) and 1(d) with $U=2.0$ and $U=3.0$, respectively. On the other hand, the area of the “intermediate phase” identified by the fidelity metric decreases but does not vanishes as $U$ increases from $0.0$ to $3.0$. We discuss the ($U$, $\Delta$) phase diagram with $V=0$ in Appendix A to better understand the influence of $U$ on the ($V$, $\Delta$) phase diagrams. We must have a statement here that in our ED discussions, the $C=1$ phase is equivalent to the “intermediate phase” only when $U=0.0$, and they are not equivalent when $U$ is finite because of the discrepancies. Similar discussions have also been reported in Refs. Shao et al. (2021); Yuan et al. (2023) and they found that it is better to use the fidelity results to characterize the phase transitions in the interacting Haldane models because the discretized method to calculate Chern number suffers from the finite-size effect more severely. The MF results in Sec. III.2 also support the existence of $C=1$ in the case of $U=3.0$. Even so, we still suggest that the question of whether the $C=1$ phase can exist with $U$ increased to $3.0$ may be open. This issue is worth further investigation, especially considering that in materials $U$ is generally larger than $V$.
@@
Figure 2: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SSO/SCO}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $V=0.4$ on the left panels and $\Delta=1.4$ on the right panels. The on-site interaction $U=0$ and the parameters are corresponding to the black dashed lines in Fig. 1(a).